### v2

@@ 3.3.3 Convergence analysis of Algorithm 1 and proof of validity under Assumption 2
This section starts with a rewording of the convergence analysis of Algorithm 1 from [2, Theorem 1], as Proposition 1. Then, Propositions 2 and 3 show that all its assumptions are all met in our context.
Proposition 1 (convergence analysis of Algorithm 1).
Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$. Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, if $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
@@
First, $\inf\Phi(\mathbb{X})=\inf\Phi(X)=\inf\varphi(\gamma(X))>{-}\infty$ holds by assumption, so $\Phi$ is bounded below. Second, let $\alpha\in\mathbb{R}$. For all $x\in\mathbb{X}$, the inclusion $x\in\Phi^{-1}({]}{-}\infty,\alpha])$ is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2. Thus $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded. Last, consider the partition of $X$ given by Assumption 2. For all $i\in\llbracket 1,N\rrbracket$$X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions. ∎
@@ 3.3.4 Proof of Theorem 2
Consider that Framework 1 is applicable and that Assumptions 1 and 2 hold. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$). Proposition 2 shows that Proposition 1 is applicable, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$), and then $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.c). Moreover, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\cap X$, Proposition 3 claims that $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$, then $\Phi$ is lower semicontinuous at $x^{*}$. In this last case, $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎
4 Application of the POf to composite optimization problems
@@
where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expression is unknown). We say that $\varphi$ is a greybox, since its explicit expression depends on some blackbox components. The POf (via the strategy formalized in Theorem 2) allows to solve such composite greybox problems, as shown below. This section solves four examples using this strategy.
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, we partition $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$. Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. Solving it is manageable with the cDSM and, according to Theorem 2, this allows to recover solutions to the large-dimensional Problem ($\mathbf{P}_{\texttt{ini}}$). Remark also that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
[…]
@@ 4.1 Mono-variable noise
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
@@
shown on Figure 3 (right). Its global minimizer is $x^{*}\triangleq 0$, with $\Phi(x^{*})=0$, and $\Phi$ is left-discontinuous at $x^{*}$. Assumptions 1 and 2 hold. Table 1 shows that the cDSM approaches $x^{*}$ closely and from the side avoiding the discontinuity. All instances return $\widetilde{x}^{*}\in[0,2E^{-10}]$. Hence they all provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 2E^{-10}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 2E^{-10}$. This empirically validates Theorem 2. We also remark that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$), even if the requirement in Theorem 2 that $\varphi$ is lower semicontinuous at all points in $\mathcal{ACC}(\gamma;x^{*})$ does not hold (at $(0,-1)\in\mathcal{ACC}(\gamma;x^{*})$).
@@
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$+9.753$ $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
$+\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
$+\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
$+e+1$ $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
$-9.753$ $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
$-\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
$-\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
$-e-1$ $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
@@ 4.2 Radial noise
shown on Figure 5 (right). Its minimizer is $x^{*}\triangleq\sqrt{2}$, with $\Phi(x^{*})=0$. Assumptions 1 and 2 hold. Table 2 shows the convergence of the cDSM towards $x^{*}$. All instances return $\widetilde{x}^{*}\in[x^{*}\pm 6E^{-11}]$, so $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})$ satisfies $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 6E^{-11}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 3E^{-5}$. This agrees with Theorem 2.
@@
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
@@ 4.3 Noise affected by a nonlinear combination of all variables
where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is assumed explicit, while $\varepsilon$ is defined via $\varepsilon(x)\triangleq\exp(\frac{1}{x-4})+\frac{\sqrt{\lvert{x-4}\rvert}}{5}$ if $x\neq 4$ and $\varepsilon(4)\triangleq+\infty$ and this information is assumed hidden. Then $\varphi$ has no global minimizer, but there exists a generalized global solution $y^{*}\approx(2.1287,1.8791)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ when $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$. The exact value of $y^{*}$ and the direction are given below. We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable: First, $y_{1}^{2}/(y_{2}^{2}+1)=1$ provides $y_{2}^{2}=y_{1}^{2}-1$ (and thus $y_{1}^{2}\geq 1$ so $y_{1}\neq 0$), and $(y_{1},y_{2})\in\mathbb{Y}(x)\cap\Omega$ raises $y_{2}=x/y_{1}$ and $y_{1}\geq 0$. Second, equalling the two expressions of $y_{2}^{2}$ provides $(y_{1}^{2})^{2}-y_{1}^{2}-x^{2}=0$, with one admissible value for $y_{1}^{2}$ as $y_{1}^{2}=\frac{1}{2}(1+\sqrt{1+4x^{2}})$ since $y_{1}^{2}\geq 1$. Hence $y_{1}$ follows (since $y\in\Omega$ forces $y_{1}\geq 0$), and then $y_{2}$ follows as well. Then, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
@@
We test eight instances of the cDSM minimizing $\Phi$. Table 3 shows their convergence towards $x^{*}$. All instances starting with $x^{0}<x^{*}$ return $\widetilde{x}^{*}\in[x^{*}-2E^{-10},x^{*}]$, which approaches $x^{*}$ closely and from the direction minimizing $\Phi$. Hence they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2. The instance starting with $x^{0}\triangleq 3\sqrt{2}\approx 4.24>x^{*}$ behaves similarly. The last three instances, which start with $x^{0}\gg 4$, converge instead to $\widetilde{x}^{*}\approx 9.27$, but this is consistent with Theorem 2 since $\Phi$ has a local minimizer around $9.27$ so $\varphi$ has a local minimizer around $\gamma(9.27)$.
@@
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$-e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
$-\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
$-\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
$+e$ $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
$+3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
$+2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$+e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
Table 3: results for eight instances of Algorithm 1 in the context of Section 4.3. In this table, $\hat{x}^{k}\triangleq x^{k}-4$. For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first $9$ digits constant in all cases.
4.4 Bidimensional noise and oracle provided by an algorithmic method

### v3

@@ 3.3.3 Proof of validity of Proposition 1 under Assumption 2
This section shows that the existing convergence analysis of Algorithm 1 (from [3, Theorem 2], and recalled in Proposition 1 above) is applicable in the context of Reformulation ($\mathbf{P}_{\texttt{r}}$). First, Proposition 2 shows that all assumptions involved in Proposition 1 are met in our context.
@@
Consider that Assumptions 1 and 2 hold. First, $\inf\Phi(\mathbb{X})=\inf\Phi(X)=\inf\varphi(\gamma(X))>{-}\infty$ holds by assumption, so $\Phi$ is bounded below. Second, let $\alpha\in\mathbb{R}$. For all $x\in\mathbb{X}$, the inclusion that $x$ lies in the sublevel set of $\Phi$ of level $\alpha$ reads as $x\in\Phi^{-1}({]}{-}\infty,\alpha])$, which is equivalent to $\Phi(x)\leq\alpha$, and then to $\varphi(\gamma(x))\leq\alpha$, and then to $\gamma(x)\in\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$, and then to $x\in\chi(\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha]))$ by Lemma 2. Thus, $\Phi^{-1}({]}{-}\infty,\alpha])$ is bounded, since by assumption $\varphi_{|\gamma(X)}^{-1}(]{-}\infty,\alpha])$ is bounded and $\chi(\mathcal{Y})$ is bounded for all $\mathcal{Y}\subseteq\mathbb{Y}$ bounded. Last, Assumption 2 yields $X=\sqcup_{i=1}^{N}X_{i}$ such that, for all $i\in\llbracket 1,N\rrbracket$, $X_{i}$ is ample and $\Phi_{|X_{i}}=(\varphi_{|\gamma(X_{i})})\circ(\gamma_{|X_{i}})$ is continuous as a composition of continuous functions. ∎
Then, Proposition 3 highlights some conditions to ensure the requirement of lower semicontinuity of $\Phi$ (involved in the last part of Proposition 1).
@@ 3.3.4 Proof of Theorem 2
Finally, we prove Theorem 2. We apply Proposition 1 to gather a solution $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), and then we derive a solution to Problem ($\mathbf{P}_{\texttt{i}}$) via Theorem 1.
Proof.
Consider Framework 1 under Assumptions 1 and 2. Let $(x^{k})_{k\in\mathbb{N}}$ be generated by Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$). Proposition 1 is applicable by Proposition 2, so all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$), so $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.c). This proves the first part of Theorem 2. If moreover $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\Phi$ is lower semicontinuous at $x^{*}$ by Proposition 3, so $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$) by Proposition 1, so $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) by Theorem 1.b). This completes the proof of Theorem 2. ∎
4 Application of the POf to composite problems
@@
where $\widetilde{\varphi}:\mathbb{Y}\times\mathbb{R}\to\mathbb{R}$ is smooth and easy to minimize and has an explicit definition, and $f:\mathbb{Y}\to\mathbb{X}$ is smooth with explicit definition (and $\dim(\mathbb{X})\in\mathbb{N}^{*}$ small), and $\sigma:\mathbb{X}\to\mathbb{R}$ and $\varepsilon:\mathbb{X}\to\overline{\mathbb{R}}$ are blackboxes (that is, their expressions are both unknown). We say that $\varphi$ is a greybox, since its explicit expression depends on blackbox components. Such composite greybox problems are compatible with the POf.
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{i}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$. Then, either Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{i}}$). Also, remark that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
This section solves four examples using the DFPOm. In Section 4.1, we solve the example from Section 1.1. In Section 4.2, $f$ returns the radius of the polar coordinates and $\gamma$ is easily tractable. In Section 4.3, $f$ returns a nonlinear combination of all variables and the expression of $\gamma$ may be calculated. In Section 4.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is solvable with a dedicated algorithmic method. In all cases, we implement the cDSM as described in Remark 5. These examples show the variety of what Theorem 2 may guarantee. In Sections 4.1 and 4.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 4.1. In Sections 4.2 and 4.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$.
@@ 4.1 Mono-variable noise
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{r}}$) minimizes
@@
[…]
@@
$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
+9.753+9.753 $\hat{x}^{23}=+3.37E^{-03}$ $\hat{x}^{37}=+8.86E^{-07}$ $\hat{x}^{55}=+3.15E^{-09}$ $\hat{x}^{64}=+7.41E^{-12}$
$\pi$ $\hat{x}^{18}=+3.12E^{-03}$ $\hat{x}^{31}=+3.93E^{-06}$ $\hat{x}^{49}=+1.00E^{-09}$ $\hat{x}^{56}=+6.89E^{-11}$
$\sqrt{2}$ $\hat{x}^{13}=+4.65E^{-03}$ $\hat{x}^{29}=+2.79E^{-06}$ $\hat{x}^{44}=+3.08E^{-09}$ $\hat{x}^{54}=+5.12E^{-11}$
+e+1+e+1 $\hat{x}^{17}=+2.08E^{-03}$ $\hat{x}^{24}=+3.01E^{-06}$ $\hat{x}^{45}=+3.29E^{-09}$ $\hat{x}^{53}=+3.45E^{-11}$
9.753-9.753 $\hat{x}^{17}=+4.61E^{-04}$ $\hat{x}^{33}=+3.33E^{-06}$ $\hat{x}^{51}=+3.96E^{-09}$ $\hat{x}^{60}=+4.05E^{-12}$
$\pi$ $\hat{x}^{20}=+2.94E^{-03}$ $\hat{x}^{33}=+3.46E^{-06}$ $\hat{x}^{43}=+2.33E^{-10}$ $\hat{x}^{54}=+1.16E^{-10}$
$\sqrt{2}$ $\hat{x}^{06}=+3.45E^{-03}$ $\hat{x}^{27}=+1.13E^{-06}$ $\hat{x}^{43}=+1.51E^{-09}$ $\hat{x}^{52}=+7.67E^{-13}$
e1-e-1 $\hat{x}^{21}=+2.74E^{-03}$ $\hat{x}^{36}=+2.40E^{-06}$ $\hat{x}^{49}=+4.02E^{-09}$ $\hat{x}^{57}=+5.70E^{-11}$
@@ 4.2 Radial noise
[…]
@@
$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
@@ 4.3 Noise affected by a nonlinear combination of all variables
where $f:y\in\mathbb{Y}\mapsto y_{1}y_{2}\in\mathbb{R}$ is explicitly known, while the blackbox $\varepsilon:\mathbb{R}\to\overline{\mathbb{R}}$ is defined via $\varepsilon(4)\triangleq+\infty$ and $\varepsilon(x)\triangleq\exp((x-4)^{-1})+\frac{1}{5}\sqrt{\lvert{x-4}\rvert}$ if $x\neq 4$. Then $\varphi$ has no global minimizer, but it has a generalized global minimizer $y^{*}\approx(2.13,1.88)$, since $\inf\varphi(\Omega)=0$ and $\varphi(y)\to 0$ as $y\to y^{*}$ from some direction, even though $\varphi(y^{*})=+\infty$. The exact value of $y^{*}$ and the direction are given below. We define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, so $\varphi_{|\mathbb{Y}(x)}$ is smooth (since $f(y)=x$ so $\varepsilon(f(y))$ is constant) and its minimizer is computable (by solving $y_{1}^{2}/(y_{2}^{2}+1)=1$ subject to $y\in\Omega$ and $f(y)=x$). Then, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is
@@
We test eight instances of the cDSM minimizing $\Phi$. Table 3 shows their convergence towards $x^{*}$. All instances starting with $x^{0}<x^{*}$ return $\widetilde{x}^{*}\in[x^{*}-2E^{-10},x^{*}]$, which approaches $x^{*}$ closely and from the direction minimizing $\Phi$. Hence, they provide $\widetilde{y}^{*}\triangleq\gamma(\widetilde{x}^{*})\approx y^{*}$ with $\varphi(\widetilde{y}^{*})\approx\inf\varphi(Y)$, which agrees with Theorem 2. The instance starting with $x^{0}\triangleq 3\sqrt{2}\approx 4.24>x^{*}$ behaves similarly. The last three instances, which start with $x^{0}\gg 4$, converge instead to $\widetilde{x}^{*}\approx 9.27$, but this is consistent with Theorem 2 since $\Phi$ has a local minimizer around $9.27$ so $\varphi$ has a local minimizer around $\gamma(9.27)$.
@@
$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-03}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-06}\right]$ $1^{st}\penalty 10000\ \hat{x}^{k}\in\left[\pm 5E^{-09}\right]$ returned $\hat{x}^{k}$
$e^{2}$ $\hat{x}^{107}=-4.17E^{-03}$ $\hat{x}^{120}=-2.20E^{-06}$ $\hat{x}^{138}=-1.68E^{-09}$ $\hat{x}^{146}=-4.65E^{-11}$
$\pi$ $\hat{x}^{026}=-1.76E^{-03}$ $\hat{x}^{043}=-1.70E^{-06}$ $\hat{x}^{057}=-3.88E^{-09}$ $\hat{x}^{065}=-4.14E^{-11}$
$\sqrt{2}$ $\hat{x}^{024}=-4.25E^{-04}$ $\hat{x}^{040}=-4.95E^{-06}$ $\hat{x}^{051}=-1.40E^{-09}$ $\hat{x}^{062}=-8.01E^{-12}$
+e+e $\hat{x}^{009}=-4.68E^{-04}$ $\hat{x}^{026}=-2.78E^{-06}$ $\hat{x}^{043}=-3.13E^{-09}$ $\hat{x}^{051}=-1.06E^{-10}$
$3\sqrt{2}$ $\hat{x}^{017}=-2.66E^{-03}$ $\hat{x}^{030}=-1.33E^{-07}$ $\hat{x}^{046}=-2.41E^{-09}$ $\hat{x}^{053}=-8.40E^{-11}$
$2e^{2}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$4\pi$ / / / $\hat{x}^{201}=+5.27E^{+00}$
$e^{3}$ / / / $\hat{x}^{201}=+5.27E^{+00}$
Table 3: Results for eight instances of Algorithm 2 in the context of Section 4.3. In this table, $\hat{x}^{k}\triangleq x^{k}-4$. For each of the last three instances, we arbitrarily return $x^{201}\approx 9.26779505$ since $(x^{k})_{k\geq 40}$ has its first 99 digits constant in all cases.
4.4 Bi-dimensional noise and oracle provided by an algorithmic method